OpenStudy (anonymous):

Can someone help me factor 5k^3+20k^2-15

OpenStudy (cwrw238):

first find the GCF of the 3 terms it must be a number because of the 15 at the end so what's the greatest common factor of 5, 15 and 20 ?

OpenStudy (anonymous):

5, right

OpenStudy (cwrw238):

yes to take 5 out and u can write: 5(k^3 + 4k^2 - 3)

OpenStudy (anonymous):

So, to factor stuff you just take out the GCF?

OpenStudy (cwrw238):

well sometimes

OpenStudy (cwrw238):

its the first thing you look for

OpenStudy (anonymous):

ok, is there anything else i need to do?

OpenStudy (cwrw238):

no - thats it

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

Well, wait a minute \[ 5 k^3+20 k^2-15=5 (k+1) \left(k^2+3 k-3\right) \]

OpenStudy (cwrw238):

oh yes - i missed that - sorry Jailee

OpenStudy (anonymous):

its fine thanks again

OpenStudy (cwrw238):

yw - these cubic expressions can sometimes be factored

OpenStudy (anonymous):

for a cubic try to find a zero by hit and trial method. try k=-1 5(-1)^3+20(-1)^2-15=-5+20-15=0 hence k=-1 or k+1 is a factor divide by -1,for synthetic division or by k+1 long division. you get other factor.